# When Hyperspectral Image Classification Meets Diffusion Models: An Unsupervised Feature Learning Framework

### v1

@@ Abstract
Learning effective spectral-spatial features is important for the hyperspectral image (HSI) classification task, but the majority of existing HSI classification methods still suffer from modeling complex spectral-spatial relations and characterizing low-level details and high-level semantics comprehensively. As a new class of record-breaking generative models, diffusion models are capable of modeling complex relations for understanding inputs well as learning both high-level and low-level visual features. Meanwhile, diffusion models can capture more abundant features by taking advantage of the extra and unique dimension of timestep $t$. In view of these, we propose an unsupervised spectral-spatial feature learning framework based on the diffusion model for HSI classification for the first time, named Diff-HSI. Specifically, we first pretrain the diffusion model with unlabeled HSI patches for unsupervised feature learning, and then exploit intermediate hierarchical features from different timesteps for classification. For better using the abundant timestep-wise features, we design a timestep-wise feature bank and a dynamic feature fusion module to construct timestep-wise features, adaptively learning informative multi-timestep representations. Finally, an ensemble of linear classifiers is applied to perform HSI classification. Extensive experiments are conducted on three public HSI datasets, and our results demonstrate that Diff-HSI outperforms state-of-the-art supervised and unsupervised methods for HSI classification.
@@
Hyperspectral image classification, denoising diffusion probabilistic model, unsupervised feature learning.
@@ I Introduction
HYPERSPECTRAL image (HSI) classification plays a crucial role in remote sensing [1, 2, 3, 4], as it aims to distinguish each pixel’s category in hyperspectral data by using dense and detailed electromagnetic spectral information [5]. The main objective is to effectively discriminate objects of interest by capturing subtle variations in the spectral signatures associated with their pixels [6]. To this end, a large number of methods are developed, relying on hand-crafted features [7, 8, 9, 10, 11]. However, these works are designed to extract features manually, leading to limited discriminative information extraction from the high spectral variability [12].
The rapid advancement of deep learning has led to the emergence of supervised approaches, a typical category of HSI classification methods, specifically utilizing deep convolutional neural networks (CNNs) [13, 14, 15] and transformers [16, 17, 18, 19]. Although these previous supervised methods have already achieved promising results in HSI classification, they still have limitations in spectral-spatial feature extraction. Firstly, previous supervised methods train the CNN or transformer network to learn explicit mappings which leads to poor performance to model complex spectral-spatial relations. Secondly, supervised methods tend to prioritize features from labeled areas while HSIs have large unlabeled regions that contain label-agnostic information. They result in a preference for learning labeling priors and limited general representation generation that characterize spectral-spatial relations. Additionally, the limited availability of labeled samples in HSI datasets poses constraints on previous supervised methods.
((a))
((b))
Figure 1: Different unsupervised feature learning frameworks including (a) existing unsupervised feature learning and (b) ours proposed Diff-HSI that extract timestep-wise features and construct multi-timestep representations by dynamic fusion for the first time.
As another category of HSI classification methods, [20, 21, 22, 23] is designed to learn spectral-spatial features from HSIs in an unsupervised manner. Typically, these methods leverage an encoder-decoder network trained in an unsupervised manner for HSI reconstruction and extract the spectral-spatial feature from the intermediate feature maps, depicted in Fig. 1(a). Although these methods mine implicit spectral attributes and spatial relations through unsupervised feature learning, the reconstruction networks they designed mainly capture low-level reconstruction features such as texture and edge, lacking high-level features which are necessary for discriminative representations to model the spectral-spatial relations comprehensively [24].
Learning rich representations and modeling the complex spectral-spatial relations from high dimensions are significant and should be considered for HSI classification. Most recently, diffusion models [25, 26, 27] have emerged as new generative models with superior performance in generation and reconstruction tasks, and have started to be explored in other computer vision tasks like semantic segmentation [28, 29, 30]. Diffusion models have the ability in the reconstruction task, so they are well-suited as a framework for unsupervised feature learning. Furthermore, there exist two appealing properties of using diffusion models [31, 32, 33] for unsupervised feature learning of hyperspectral data: 1) diffusion models learn implicit mappings, which have superior representation ability on complex problems via simple expressions and extract both high-level and low-level features from images; 2) the stepwise reverse denoising process of diffusion models are modeled as an iterative optimization process optimized by Langevin dynamics. This process brings a large number of degrees of freedom and tries to predict further information conditioned on the given noise-destructed data at every timestep, leading to better generalization to complex problems. When applying diffusion models to unsupervised feature learning of HSI data, it not only learns both high-level and low-level features but also extracts spectral-spatial features in the dimension of timestep $t$, as shown in Fig. 1(b).
Naturally, it is crucial to explore how to exploit abundant timestep-wise features more effectively and efficiently. On the one hand, the timestep-wise features are diverse and informative and may focus on different information. The shallow timestep-wise features are more informative for fine details, while deeper ones are more concerned about high-level semantics and global information [25, 34]. On the other hand, HSIs from various datasets acquired by different sensors exhibit distinct spectral characteristics, leading to variations in the spectral representation patterns of timestep-wise features extracted from different HSI datasets. Meanwhile, the abundant timestep-wise features pose a challenge in terms of computation overheads. In addition, manually selecting timestep-wise features to characterize spectral-spatial features effectively is struggle. Therefore, a dynamic feature fusion module that adaptively fuses features of different timesteps for different HSI datasets is desired.
@@
In view of these, we propose the unsupervised spectral-spatial feature learning framework based on the diffusion model for HSI classification, named Diff-HSI. More specifically, the proposed framework first pretrain a diffusion model with unlabeled HSI patches for unsupervised feature learning, which is able to automatically learn label-agnostic knowledge and mine the connotation of unlabeled data that reveals the complex spectral-spatial dependencies compared to conventional supervised methods. Then, we extract features from different hierarchies from the pretrained denoising U-Net decoder to construct the timestep-wise feature bank, which contains various informative features from different timestep $t$. To effectively harness the features from the timestep-wise feature bank and softly learn the proper timestep-wise feature combination for different HSI datasets, we propose a dynamic feature fusion module. This module is designed to adaptively fuse hierarchical features from different timesteps, generating multi-timestep representations enriched with comprehensive spectral-spatial information. Ultimately, an ensemble of linear classifiers is employed to leverage these dynamic representations for accurate HSI classification.
@@
    As a pioneer work, we introduce diffusion models to HSI classification and propose a diffusion-based unsupervised feature learning framework called Diff-HSI, which can implicitly learn both high-level and low-level features to model complex spectral-spatial relations.
@@
    To leverage abundant timestep-wise features effectively and efficiently, we construct a timestep-wise feature bank with hierarchical features from different timesteps, and design a dynamic feature fusion module to fuse these timestep-wise features adaptively. The above process learns the soft timestep-wise feature combination for each labeled patch of different HSI datasets and constructs comprehensive multi-timestep representations.
@@
    Compared with several state-of-the-art supervised and unsupervised methods, experimental results demonstrate that our proposed method achieves superior classification accuracy on three public HSI datasets.
The remainder of this paper is organized as follows. Section II describes related work. In Section III, our proposed Diff-HSI is introduced in detail. Section IV conducts extensive experiments on three HSI datasets to demonstrate the effectiveness of the proposed method. Finally, some conclusions are drawn in Section V.
@@ II Related Work
II-A Supervised-based HSI Classification Methods
Due to the remarkable breakthroughs achieved by deep learning in various computer vision tasks, many progressive deep learning-based networks have been widely utilized for supervised HSI classification methods, e.g., autoencoders (AEs) [21], deep belief networks (DBNs) [35] and recurrent neural networks (RNNs) [36]. Among these, CNNs draw significant attention with their excellent feature extraction capability and become mainstream in HSI classification [13, 14, 15]. CNNs can effectively extract spectral-spatial features through their ability to extract spatially structural information and locally contextual information. As the vision transformer rises in the field of computer vision, researchers begin to explore the value of the transformer in HSI classification and find it outperforms CNN-based methods in solving the problem of long-term spectral information dependencies [16, 17, 18, 19]. Although these supervised methods have achieved promising performance, they still require large amounts of labeled HSI samples, and are limited to modeling complex relations and capturing label-agnostic information.
II-B Unsupervised-based HSI Classification Methods
@@
Unsupervised feature learning is a feature extraction paradigm to automatically learn feature representations from the input data without any annotated information. Leveraging unsupervised feature learning on unlabeled HSI patches is a solution to the limited labeled samples of HSI datasets. Typically, the commonly used unsupervised feature learning methods in HSI classification are based on the encoder–decoder paradigm, where an autoencoder-like network encodes the input HSI patches into a refined feature and then reconstructs the feature to initial HSI data by a decoder. Mou et al. [20] first design a fully 2D Conv–Deconv network in an end-to-end manner for unsupervised feature learning of HSI classification. Similarly, Mei et al. [21] design a 3D convolutional autoencoder (3D-CAE) for unsupervised feature learning of HSI classification. To alleviate the insufficiency of geometric representation and exploit the multi-scale features, Zhang et al. [22] design a multi-scale CNN-based unsupervised feature learning framework, with two branches of decoder and clustering optimized by the error feedback of image reconstruction and pseudo-label classification. In the aforementioned approaches, the feature representations learned from the unlabeled HSI patches contain more spectral-spatial information than supervised methods; however, they mainly focus on low-level visual concepts and ignore high-level feature extraction for HSI classification, which is solved in our methods.
II-C Diffusion Models
@@
Diffusion models are a class of probabilistic generative models that progressively inject a standard Gaussian noise, then learn a model to reverse this process for sample generation [25, 26, 27]. Current research on diffusion models is mostly based on three formulations: denoising diffusion probabilistic models (DDPMs) [25], score-based generative models [33], and stochastic differential equations [37]. DDPMs are the mainstream diffusion models, and a large number of recent works based on DDPMs have made DDPMs increasingly powerful in terms of generative quality and diversity over other generative models [38]. Meanwhile, DDPMs have been widely used in several applications, including super-resolution [39], inpainting [40], and point cloud generation [41]. However, to the best of our knowledge, few works successfully adopt them to HSI classification. When applying diffusion models to HSI classification, it can not only implicitly learn both high-level and low-level features but also extract spectral-spatial features in the dimension of timestep $t$, leading to improvement in modeling complex spectral-spatial relations.
Figure 2: Overview of our proposed Diff-HSI. The method consists of two steps. Step 1: We pretrain the DDPM with HSI patches in an unsupervised manner and then freeze the parameters. Step 2: We construct a timestep-wise feature bank by extracting hierarchical features from the fixed pretrained U-Net decoder in terms of different timestep $t$. In order to softly fuse the features, a dynamic feature fusion module is designed to obtain effective multi-timestep representations adaptively. Finally, an ensemble of lightweight classifiers is trained by leveraging the multi-timestep representations of label-limited HSI patches to perform HSI classification.

### v2

# Exploring Multi-Timestep Multi-Stage Diffusion Features for Hyperspectral Image Classification

@@ Abstract
The effectiveness of spectral-spatial feature learning is crucial for the hyperspectral image (HSI) classification task. Diffusion models, as a new class of groundbreaking generative models, have the ability to learn both contextual semantics and textual details from the distinct timestep dimension, enabling the modeling of complex spectral-spatial relations in HSIs. However, existing diffusion-based HSI classification methods only utilize manually selected single-timestep single-stage features, limiting the full exploration and exploitation of rich contextual semantics and textual information hidden in the diffusion model. To address this issue, we propose a novel diffusion-based feature learning framework that explores Multi-Timestep Multi-Stage Diffusion features for HSI classification for the first time, called MTMSD. Specifically, the diffusion model is first pretrained with unlabeled HSI patches to mine the connotation of unlabeled data, and then is used to extract the multi-timestep multi-stage diffusion features. To effectively and efficiently leverage multi-timestep multi-stage features, two strategies are further developed. One strategy is class & timestep-oriented multi-stage feature purification module with the inter-class and inter-timestep prior for reducing the redundancy of multi-stage features and alleviating memory constraints. The other one is selective timestep feature fusion module with the guidance of global features to adaptively select different timestep features for integrating texture and semantics. Both strategies facilitate the generality and adaptability of the MTMSD framework for diverse patterns of different HSI data. Extensive experiments are conducted on four public HSI datasets, and the results demonstrate that our method outperforms state-of-the-art methods for HSI classification, especially on the challenging Houston 2018 dataset. The codes are available at https://github.com/zjyaccount/MTMSD.
@@
Hyperspectral image classification, denoising diffusion probabilistic model, multi-timestep multi-stage features, feature purification, feature selection.
@@ I Introduction
HYPERSPECTRAL image (HSI) classification plays a crucial role in remote sensing, as it aims to distinguish each pixel’s category in hyperspectral data by using dense and detailed electromagnetic spectral information [1, 2]. HSI classification has a wide range of applications in environmental monitoring [3], resource management [4], agriculture disaster response [5], military defense [6], etc.
The extraction of spectral-spatial features from HSIs plays a significant role in HSI classification tasks. Initially, machine learning-based feature extraction methods have been developed [7, 8, 9, 10, 11]. However, these approaches rely on manual feature engineering, resulting in limited extraction of discriminative information from the highly variable spectral data [12]. In recent years, the rapid advancement of deep learning has paved the way for neural network-based approaches for feature extraction, specifically leveraging convolutional networks (CNNs) [13, 14, 15] and transformer models [16, 17, 18, 19]. These neural network-based methods excel at automatically learning valuable spectral-spatial features from labeled HSI data, thus achieving promising results in HSI classification. Further, unsupervised methods such as [20, 21, 22] have been designed to excavate spectral-spatial features from unlabeled HSIs. Typically, these methods employ an encoder-decoder network trained in an unsupervised manner for HSI reconstruction, thus extracting spectral-spatial features from unlabeled HSI data. Leveraging unsupervised feature learning enables deep mining of large amounts of unlabeled regions in HSI, which contain a wealth of label-agnostic information, thereby promoting the HSI classification task.
Figure 1: Overview of existing diffusion-based feature learning frameworks for the HSI classification task. Our method can fully explore and exploit rich contextual semantics and textual features hidden in the diffusion model.
Recently, diffusion models [23, 24, 25] have emerged as powerful models with superior performance in generation and reconstruction tasks. These models have also been explored in various computer vision tasks such as semantic segmentation [26, 27, 28]. Distinguished from traditional deep neural networks, diffusion models employ a stepwise reverse denoising process, formulated as an iterative optimization procedure optimized by Langevin dynamics [29]. This approach introduces a multitude of degrees of freedom for feature learning and predicts additional information conditioned on the given noise-corrupted data at each timestep. Consequently, diffusion models can implicitly capture both high-level and low-level visual concepts, facilitating better generalization and modeling of complex spectral-spatial relations [30, 31]. Therefore, one recent work has proposed using diffusion models for HSI classification [32], which employs unsupervised feature learning to extract diffusion features. However, the diffusion features used in [32] are extracted solely from a single timestep and a single stage of the denoising U-Net, and these selections about the timestep and stage are manually determined based on extensive experimentation with each dataset, as shown in Fig. 1. Firstly, relying on single-timestep features from a single U-Net layer inevitably results in the loss of abundant spectral-spatial information and limits the effectiveness of modeling spectral-spatial relations. Secondly, the manual selection approach lacks generality and adaptability to diverse HSI datasets exhibiting specific spectral representation patterns.
Naturally, it is crucial to explore how to exploit abundant multi-timestep multi-stage features extracted from the whole stage of denoising U-Net more effectively and efficiently. First, multi-timestep features extracted from diffusion models are diverse and focus on different information. The shallow timestep-wise features are more informative for textual details, while deeper ones are more concerned about high-level semantics and global information [23, 33]. For modeling complex spectral-spatial relations, textual features depict the spatial distribution and changing patterns of ground objects, and contextual semantics represent the spectral attributes and the content of ground objects. Thus, leveraging multi-timestep features can integrate both contextual semantics and textual features to build comprehensive spectral-spatial representation for better performance. Second, multi-stage features, as compared to single-stage features, encapsulate a rich hierarchy of information that contains richer semantic and reconstruction characteristics, facilitating the modeling of spectral-spatial relations. Moreover, [32] also demonstrates that various stage features improve classification performance with different degrees across various datasets. Although multi-timestep multi-stage features are desired, HSIs from various datasets acquired by different sensors exhibit distinct spectral-spatial characteristics, leading to variations in the spectral-spatial representations. Moreover, for different regions in the same HSI, the emphasis on texture and semantics varies, leading to distinct preferences in the selection of timestep $t$. Meanwhile, multi-stage features are numerous and comprise both redundant semantics and reconstruction information along the channel dimension, posing a challenge in terms of memory constraints when training and inferring on large datasets.
@@
In view of these, we propose a novel diffusion-based feature learning framework that explores Multi-Timestep Multi-Stage Diffusion features for HSI classification for the first time, named MTMSD. More specifically, the proposed framework first pretrain a diffusion model with unlabeled HSI patches for diffusion feature learning, mining the connotation of unlabeled data that reveals the complex spectral-spatial dependencies. Then, we extract multi-timestep multi-stage diffusion features from the pretrained denoising U-Net decoder and construct the timestep-wise center and global feature bank by center extraction and average pooling. After that, two strategies are further developed in MTMSD to leverage multi-timestep multi-stage diffusion features effectively and efficiently. First, to reduce the redundancy of multi-stage features and maintain efficiency, we propose to perform class & timestep-oriented multi-stage feature purification on multi-stage features in the timestep-wise center feature bank with the inter-class and inter-timestep prior. Second, to effectively harness multi-timestep features and softly learn the proper timestep-wise feature combination for different datasets, we propose the selective timestep feature fusion module. This module is designed to adaptively select different timestep center features with the guidance of related global features, and fuse them for multi-timestep multi-stage selective representations that integrate contextual semantics and textual features to model comprehensive spectral-spatial relations. Ultimately, an ensemble of linear classifiers is employed for accurate HSI classification.
@@
    For modeling complex spectral-spatial relations, we propose a novel diffusion-based framework that explores multi-timestep multi-stage diffusion features for HSI classification. To the best of our knowledge, this is the first work to learn and exploit multi-timestep multi-stage diffusion features for diffusion-based HSI classification.
@@
    We design the novel class & timestep-oriented multi-stage feature purification module. It adaptively selects significant channels of multi-stage diffusion features from both inter-class and inter-timestep aspects to reduce redundant information and maintain computational efficiency.
@@
    We propose to perform selective timestep feature fusion on multi-timestep diffusion features. This module allows each labeled patch of different datasets to adaptively select different timestep center features with the guidance of related global features, and fuse them for comprehensive multi-timestep multi-stage selective representations that integrate contextual semantics and textual information.
  • 4)
    Compared with several state-of-the-art HSI classification methods, experimental results demonstrate that our proposed method achieves significant classification accuracy on four public HSI datasets, especially on the challenging Houston 2018 dataset.
The remainder of this paper is organized as follows. Section II describes related work. In Section III, our proposed MTMSD is introduced in detail. Section IV conducts extensive experiments on four HSI datasets to demonstrate the effectiveness of the proposed method. Finally, some conclusions are drawn in Section V.
@@ II Related Work
II-A HSI Classification
HSI classification is an important research topic in the area of remote sensing. Since HSI classification aims to distinguish each pixel’s category in hyperspectral data using dense electromagnetic spectral information [1], a large number of handcrafted feature-based methods have been designed for HSI classification [7, 8, 9, 10, 11]. Several works adopt morphological profiles (MPs) for manually extracting spectral-spatial features from HSIs. They achieve good classification results using MPs as input vectors with a support vector machine classifier. Subspace-based learning, such as sparse representation and manifold learning, is another common feature extraction strategy for HSI classification. These methods transform the high-dimensional original space using a low-dimensional subspace representation to learn spectral-spatial information.
Due to the remarkable breakthroughs achieved by deep learning in various computer vision tasks, many progressive deep learning-based networks have been widely utilized for HSI classification methods. Among these, CNNs draw significant attention with their feature extraction capability to extract spatially structural information and locally contextual information and become mainstream in HSI classification [13, 14, 15]. Based on the spectral and spatial attention modules, Zhu et al. [14] embed a residual block into a sequential spectral-spatial feature learning network. This architecture not only mitigates the risk of overfitting but also enhances classification performance. However, CNNs have the challenge of modeling long-term spectral information dependencies. To address this defect, researchers explore the value of the transformer and widely leverage it in HSI classification [16, 17, 18, 19]. Hong et al. [16] consider the spectral sequence of neighboring bands and design a pure transformer-based SpectralFormer (SF) backbone network, representing sequence attributes of spectral signatures. Sun et al. [17] extract high-level semantic features by introducing a Gaussian weighted token module into the transformer architecture, achieving promising performance in both classification accuracy and computational complexity.
@@ II-A HSI Classification
In addition to supervised feature learning, unsupervised feature learning aims to learn feature representations from the input data without any annotated information, providing a solution to the limited labeled samples of HSI datasets. Typically, the commonly used unsupervised feature learning methods in HSI classification are based on the encoder–decoder paradigm, where an autoencoder-like network encodes the input HSI patches into a purified feature and then reconstructs the feature to initial HSI data by a decoder network. Mou et al. [34] first design a fully 2D Conv–Deconv network in an end-to-end manner for unsupervised feature learning of HSI classification. Similarly, Mei et al. [20] design a 3D convolutional autoencoder (3D-CAE) for unsupervised feature learning of HSI classification. To alleviate the insufficiency of geometric representation and exploit the multi-scale features, Zhang et al. [21] design a multi-scale CNN-based unsupervised feature learning framework, with two branches of decoder and clustering optimized by the error feedback of image reconstruction and pseudo-label classification.
More recently, with the rise of diffusion models, one recent work propose a diffusion-based HSI classification method. Chen et al. propose SpectralDiff [32] that extracts the spectral-spatial diffusion features from spectral–spatial denoising network and directly feeds them into the attention-based classification network for classification. However, they only use a single timestep feature from a single stage of the denoising U-Net, which is manually selected according to extensive experiments on each dataset. This results in a lack of information to model spectral-spatial relations and poor robustness to diverse HSI datasets with diverse spectral characteristics. In our work, we propose a novel diffusion-based HSI classification framework that explores multi-timestep multi-stage diffusion features. Through multi-stage feature purification and selective timestep fusion, our MTMSD enables adaptive integration of both contextual semantics and textual features to model complex spectral-spatial relations and generality for diverse patterns of different HSI data.
II-B Diffusion Models
@@
Diffusion models are a class of probabilistic generative models that progressively inject a standard Gaussian noise, then learn a model to reverse this process for sample generation [23, 24, 25]. Current research on diffusion models is mostly based on three formulations: denoising diffusion probabilistic models (DDPMs) [23], score-based generative models [31], and stochastic differential equations [35]. Among them, DDPMs are the mainstream diffusion models, and a large number of recent works based on DDPMs have made DDPMs increasingly powerful in terms of generative quality and diversity over other generative models [36]. Meanwhile, DDPMs have been widely used in several applications, including super-resolution [37], inpainting [38], and point cloud generation [39]. Recently, [40] proposes a simple diffusion-based semantic segmentation approach that exploits 3-timestep multi-stage features with manually selected timesteps and also proves that diffusion features capture high-level semantic information for semantic segmentation. However, a fixed timestep set results in suboptimal and non-generic features for different datasets. Additionally, such a simple utilization of diffusion features for segmentation may lead to inefficiencies due to redundant information and ineffectiveness in building semantic representations from diffusion features. Differently, we focus on the HSI classification task, and dynamically select the timesteps and stages of multi-timestep multi-stage features ranging from low to high, integrating textural features and semantics to build comprehensive spectral-spatial representations.
II-C Deep Feature Selection
Feature selection [41], an essential process in deep-learning-based computer vision, plays a pivotal role in improving model performance. The goal of feature selection is to retain informative and refined features from the original features, thereby reducing dimensionality and computational complexity. For different images, the important features vary, and feature selection assists models in adapting to data from diverse domains, extracting significant and refined features. Early feature selection methods select features with input-dependent soft attention. Hu et al. [42] introduce a channel attention module named SENet to adaptively recalibrate channel-wise features by exploiting the inter-channel relations. Similar to channel attention, spatial attention methods GENet [43] is designed to enhance a network’s capacity for context information modeling via spatial masks. Building on these, Woo et al. [44] combine both channel and spatial attention, introducing the CBAM method. Additionally, feature selection on different kernel features is also a self-adaptive and effective mechanism. Li et al. [45] propose SKNet to select the features extracted from different convolutional kernels using softmax attention along the channel dimension. Different from the above methods, our framework is the first to use feature selection on multi-timestep multi-stage features for HSI classification. In detail, MTMSD creatively employs inter-class and inter-timestep priors for multi-stage feature purification in the channel dimension, and concurrently conducts feature selection on multi-timestep features with the guidance of global information.
Figure 2: Overview of our proposed MTMSD. The method consists of two steps. Step 1: We pretrain the DDPM with HSI patches in an unsupervised manner for diffusion feature learning. Step 2: We extract multi-timestep multi-stage diffusion features from the pretrained denoising U-Net decoder and construct the timestep-wise center and global feature bank by center extraction and average pooling. To effectively and efficiently leverage multi-timestep multi-stage features, we first perform class & timestep-oriented multi-stage purification on multi-stage features in the timestep-wise center feature bank, and then, we perform selective timestep feature fusion with global-feature guidance on the purified timestep-wise center feature bank. Classification is performed through an ensemble of lightweight classifiers.